# Embodying Control in Soft Multistable Grippers from Morphofunctional Co-design

### v2

@@ 2.2 Model derivation and mechanics-informed parameter determination
A comparison between FE simulations and the proposed model for different stable states of the finger (see Figure 3b) and different design cases (see Figure S10) shows good agreement with the energy-based lattice model, while reducing the computational time by more than $10^{-3}\times$ (see Table S2). The model is further extended to capture the structure’s dynamic response, allowing us to observe the material’s viscoelastic behavior and analyze metastable cases that energy minimization alone cannot predict. The dynamic response for different unit cell geometries is shown in Figure S7b, and for different five-segmented fingers in Figure S7d, where the dynamic model predicts the same final state of the system after reaching a steady state. To explore the combined behavior of different dome geometries, we modeled a five-segment DPF topology with bistable and metastable units (see Figure 3c). We stacked four bistable units ($H_{1}=H_{2}=H_{3}=H_{4}=5$ mm) and two metastable units with the same dome height ($H_{5}=H_{6}=3$ mm). This configuration allows us to achieve one unique set point: a stable state and a time-dependent state that holds its shape for a given predictable time $\tau$ (see Figure 3c). The resulting dynamic response features five different snap-through events and two snap-backs (dome-unit resets), adequately representing the complex dynamics displayed by the DPF (Figure 3c).
2.3 Position and stiffness inverse design
@@
Based on the model’s performance and its agreement with the FE simulations, we pose an inverse problem to co-design different DPF topologies capable of reaching a desired pre-programmed final position to achieve embodied position control. Since the stable positions of the system are already determined by the stable states, the optimization algorithm only searches within the fully actuated configurations, making it computationally efficient. We utilized Bayesian optimization [51] to iterate over various numbers of segments, dome heights for each of the segments ($H_{i}$), thickness of the dome ($t$), unit separation for each of the segments ($\text{U}^{i}_{\text{sep}}$), and unit length for each of the segments ($\text{U}^{i}_{\text{L}}$) (see Figure 2a for parameter reference). Two different objective functions are utilized: the first based on position and targeting a specific tip coordinate (Position Design), and the second one targeting a specific tip coordinate while maximizing the finger’s stiffness (Position + Stiffness Design; see section A.6).
@@
Position Design: Given a target tip coordinate [x,y] ($\text{Target}_{xy}$), the inverse problem objective function can be written as:
$\min_{H_{i},t,\text{U}^{i}_{L},\text{U}^{i}_{\text{sep}}}\left(\text{Target}_{xy}-\text{Tip}_{\text{dis}}\right)^{2}$ (6)
@@
where the $\text{Tip}_{\text{dis}}$ represents the tip point (see Tip in Figure 2a), which coincides with the top node of the finger’s last unit. A domain constraint is imposed to guarantee that every dome unit is higher than the one in front ($H_{i+1}\leq H_{i}$), which in certain cases can lead to up to N possible stable states (N been the number of segments). Moreover, we impose constraints for all unit separations and unit lengths to be equal for simplicity and design domain to guarantee bistability on all units, based on the phase diagram shown in Figure S1b. However, different restrictions can be imposed depending on the DPF design purpose, and parameters can be individually tuned to achieve a large variety of positions and stable states (see Movie 3).
Our algorithm was tested by defining five different tip coordinate targets using two materials with distinct properties (NinjaFlex 85A and Cheetah 95A) optimized to achieve the desired positions. For each case, the target coordinates in $x$ and $y$ were provided to the algorithm (Target Position in Figure 4a), and the optimization was performed by incrementally adding one unit per run and calculating the objective function (Equation 6). To reduce the computational cost, a maximum of 8 units was specified (Figure 4b), although additional units can be included if the objective requires a larger finger geometry. The geometry with the minimum objective function value (Obj in Tables S3 and S4) was selected and plotted to verify whether the desired position was achieved (see Figure 4a).
Geometric parameters, coordinate targets, and objective function values are reported in Table S3 and S4. The final stable state and its comparison with each target can be observed in Figure S11. To validate our optimization algorithm, we 3D printed the five different geometries shown in Table S9 using both tested materials (NinjaFlex 85A and Cheetah 95A) and measured the final tip position after commanded dome inversion. An average error of 7.3% was measured across all samples, showing good agreement with the optimization algorithm and the lattice model.
@@
The topologies generated by this methodology exhibit deterministic behavior despite the use of polymeric materials, meaning that kinematic states are consistently attained as long as the gripper’s geometric parameters remain unchanged (see Figure 4c and Movie 4). Specifically, we cycle each gripper topology between the initial and fully activated states, verifying the gripper’s ability to achieve the same aperture and positioning across multiple cycles (Tests 1-3 in Figure 4c). Furthermore, we assessed material and time degradation effects by performing 300 cycles over three weeks (see Figure S17), which demonstrates that even after numerous activations and an extended period, the gripper’s performance remains within 5% of the initial design behavior. This characteristic enables robust robot design, as the bistable elements provide intrinsic kinematic control dictated by the gripper’s mechanical response. Notice that the pressure for activating the units is not controlled using sensors and closed-loop control, illustrating the embodied self-regulation from the designed multistability. Consequently, the position control inherent to the designed multistable structure serves as control set points that can be embodied in the robot’s morphology and achieved with a single, open-loop controlled actuator.
Figure 4: DPF inverse design and experimental validation. (a) Predicted results for Target 1 and Target 3 from the optimization model (Position Design). (b) Objective function values as a function of the number of segments, with the minimum value identified for each target position. (c) The gripper achieves multiple stable kinematical states predicted by the energy-based model, which are successfully transferred to 3D-printed prototypes. (d) Stiffness determination of the DPF using a follower force approach, with Normalized strain energy vs Displacement curves shown for Objectives 1, 3, and 5. (e) Comparison of the proposed model and experimental results for position-only (Pos) optimization and combined position + stiffness optimization (Stf).
Position + stiffness design: The finger’s stiffness is calculated by using the dynamic model described in section A.4, by perturbing the structure from a programmed stable configuration (i.e., set point). We determine the system’s total internal force by solving the dynamic equation:
$[M]\ddot{x_{i}}+F_{d}^{int}(x_{ij},\dot{x_{i}},\dot{x_{j}})+F_{d}^{iso}(\dot{x_{i}},\dot{x_{j}})+F_{\text{in}}=F_{ext}(x_{i,j},t)$ (7)
$F_{\text{in}}=\nabla\left(E_{L}(x_{ij})+E_{NL}(x_{ij})+E_{T}(x_{ij})\right)$
where $F_{\text{ext}}(x_{i,j},t)$ is a follower force applied at the tip of the DPF (see Figure 4) to perturb the DPF from its equilibrium position. The gradient of the internal forces in the structure, $F_{\text{in}}$, is obtained through symbolic differentiation of the expression described in Section 2.2. The force is applied at a constant load rate of 5 mm/min, ensuring a quasi-static regime. The stiffness of the DPF in the desired state is calculated by determining the internal force ($F_{\text{in}}$) and the corresponding tip displacement over time (see Figure 4). A linear fit is then applied to the resulting data. Using this methodology, the DPF’s geometry can be optimized to achieve a specific target position with maximum stiffness in the vicinity of that stable state (see Figure 4e). Equation 6 can thus be reformulated as a weighted optimization problem expressed as:
$\min_{H_{i},t,\text{Unit}^{i}_{L},\text{Unit}^{i}_{s}}w_{1}\left(\text{Target}_{xy}-\text{Tip}_{\text{ dis}}\right)^{2}+w_{2}\left(\frac{K_{\text{dis}}}{dF_{\text{in}}/dx}\right)^{2}$ (8)
@@
where $w_{1}$ and $w_{2}$ are the weights assigned to each objective, and $K_{\text{dis}}$ is the stiffness obtained from position inverse design (Equation 6). The number of segments (N) is kept consistent with the results obtained from the previous optimization (Equation 6), meaning that $H_{i}$, $\text{Unit}^{i}_{L}$, $t_{\text{lim}}$ and $\text{Unit}^{i}_{s}$ are considered to maximize stiffness. The results of the stiffness and position optimization are shown in Table S6 for Cheetah 95A and in Table S5 for Ninjaflex 85A. Moreover, we experimentally test Target 5 geometries produced by both optimization procedures (see Figure S16 and Movie 5). By following this approach, we had increased the DPF initial stiffness on average 1.94 times for all cases, achieving maximum stiffnesses of 0.55 N/mm, 0.3 N/mm, and 0.15 N/mm for cases 2, 4, and 5 respectably (see Figure 4e for case 5), which consequently enhances the achievable grasping force.
2.4 Dome Phalanx Robot (DPR) performance
To explore the capabilities of our design approach and the advantages of leveraging multistability in soft robotics, we fabricated a soft robot (Dome Phalanx Robot - DPR) by combining four DPFs ($A_{1}$ to $A_{4}$), as shown in Figure 5a. We assembled the four DPFs using 3D printing a rigid base that prevents interaction between the fingers, imposes appropriate boundary conditions our model considers, and couples our robot to a UR3 robotic arm (see Movie 6). This robot can be assembled by either combining fingers with the same topology ($A_{1}=A_{2}=A_{3}=A_{4}$), resulting in uniform global functionality, or by using two sets of actuators ($A_{1}=A_{3}$ and $A_{2}=A_{4}$), which provides multiple functionalities within the same robot body.
We demonstrate different operation modes using the proposed design methodology and FDM 3D printing to create multiple robot topologies. Specifically, we focus on three operation modes: 1) A multistable robot with at least three set points/stable states (see Figure 5a); 2) A configuration with two finger pairs that can exert different grasping forces but reach the same final position (see Figure 5b); and 3) A combination of bistable and metastable dome geometries (see Figure 2c and d), where the DPR’s constitutive material viscoelastic response of the material allows us to embody a pick and place robotic behavior into the robot geometry (see Figure 5c and d and Movie 7). For each of these operation modes, different constraints (detailed below) are imposed on the model to ensure that the structure is accurately designed for each specific task.
Figure 5: Dome Phalanx Robot (DPR) assembly and modes: Fingers 1 to 4 ($A_{1}$ $A_{4}$) can be changed for different operation modes and mechanical behaviors. a) Multistable robot with four different set points: (i) Stable state 1 with 0 inverted units, (ii) Stable state 2 with two inverted units, (iii) Stable state 3 with four inverted units, and (iv) Stable state 4 with six inverted units. b) $A_{1}$ and $A_{3}$ are optimized for position and stiffness, while $A_{2}$ and $A_{4}$ are optimized for position. $A_{2}$ and $A_{4}$ exert $F_{Dis}$, and $A_{1}$ and $A_{3}$ exert $F_{\text{stt}}$ ($F_{\text{stt}}>F_{\text{Dis}}$). c) Bistable + Metastable finger behavior: Strain energy and pressure vs. time response. d) Snapshots of embodied pick-and-place tasks: A single pressure input is given for the robot to pick the object (Commanded response) and then release it depending on the viscoelastic material response (Programmed response)(see Movie 7).
1) Multistable soft robot: We further leverage the characteristics of our multistable DPF by programming different set points for the same robot morphology (i.e., DPFs geometries). Using the optimization algorithm, we can program three stable states by minimizing for three discrete configurations (2, 4, and 6 active units). This design approach can be implemented by modifying the objective function as follows:
$\min_{H_{i},t,\text{Unit}^{i}_{L},\text{Unit}^{i}_{s}}w_{j}\left(\text{Target j}_{xy}-\text{Tip}_{\text{dis j}}\right)^{2}+w_{M+1}\left(\frac{K_{\text{dis}}}{dF_{\text{in}}/dx}\right)^{2}$ (9)
[…]
where $M$ is the number of targeted stable states, and $j$ is the stable state number ($0<j\leq M$). By doing this, we can optimize for three target positions and maximum stiffness in the fully activated state (six inverted units). For this design, we impose the number of set points, the target tip position for each set point, and the number of units. It is worth noting that for this case, there is a different constraint for the dome height (i.e., $H_{i+1}>H_{i}$), which allows each of the states to be useful for grasping objects while being accessible in a sequential manner. While the model could be expanded to modify the unit cell’s dimensions further and determine the optimal number of units, for simplicity, we fixed the number of units at six to ensure at least two active domes per stable state. We tested our system by performing an inverse design with three different target positions (see section A.6.3 for reference), using the same methodology described previously (section 2.3). An average error of 12% was found between the desired positions and the three final stable states, as shown in Figure 5a ((i) - (iv)).
This methodology demonstrates the versatility of our DPF and provides an alternative approach to encoding multiple set points within the finger’s body by modulating its unit cell geometry. The local geometric modifications create more accessible stable states, which can be controlled using a single, open-loop driven pressure inlet per DPF. Since the systems’s curvature and stiffness are functions of the number of active units (see Figure S15b), the grasping force and carrying load capacity of the DPR also become functions of the active domes. Consequently, the robot can also be programmed with different set points for force (see Movie 6). Moreover, to showcase the gripper’s versatility to grasp various objects and its carrying capacity, we performed grasping of different objects on different stable states of the DFR (see Figure S18 and Movie 6). Notably, we observed that the restriction of the object can reset some of the dome units, which makes the DFP adapt to the object shape and perform a better grip depending on the object size. A maximum load-to-weight ratio of 10.25 is determined by loading the robot until it can not withstand the weight (see Movie 6).
2) Force-adaptable robot: For the second operation mode, we utilized the results shown in section 2.3, where two different objective functions were considered to create two different finger topologies that achieve the same tip position but exert different grasping forces. Using the data from Table S3 and Table S5, we designed a robot arrangement where two fingers ($A_{1}$ and $A_{3}$) produce more grasping force than the others ($A_{2}$ and $A_{4}$). We tested the grasping force using a compliant mechanism composed of two bars that compress as the grasping force increases (see Figure 5b). We characterized the mechanism’s force vs. aperture to measure the robot’s grasping force (see Movie 5). Our experiments show that the two different objective functions result in a force increase of 1.8 times ($F_{{\text{Dis}}}$ and $F_{{\text{stt}}}$) while reaching the same position in their fully activated stable state.
@@ 2.4 Dome Phalanx Robot (DPR) performance
3) Embodied soft robotic tasks: In the third operation mode, we leverage our ability to program our system’s dynamics. By leveraging our dome structures’ viscoelastic response and geometry, we can pre-program the duration for which the structure remains activated before returning to its original state. As demonstrated in section 2.1, we can modulate the geometry to create a fully metastable finger (see Figure 2d and e). However, the same finger can exhibit bistable and metastable behaviors by combining bistable and metastable dome units. This approach allows for encoding a desired set of points into the finger and operate in the vicinity of that point by actuating the metastable unit cells (See Movie 2). Combining the meta- and bi-stability of the units enables us to design task planning operations (such as pick-and-place; see Movie 7), where the robot can pick an object by activating all units, move to a specific position, and release the object without additional interventions and with just one pressure input (i.e., eliminating the need for a release command; see Figure 5d). We tested this operation mode by designing a six-segment finger with four bistable and two metastable units. The set point with four active units is designed with $H_{1}=4$ mm, $H_{2}=4.5$ mm,$H_{3}=5$, and $H_{4}=5$ mm, while the metastable units are designed to maximize the resetting time for two different loading times (section A.6.4). We approximate the loading profile, by measuring the pressure experimentally and approximating it as shown in Figure S13. The obtained results show how an object can be held while all units are inverted, even when the system releases pressure (Figure 5c). However, the object is automatically released from the gripper after the metastable units reset, as desired from the programmed dynamics. Using this approach, we can program a simple pick-and-place operation by activating all domes or perform a simple pinch operation as demonstrated in Figure 5d and Movie 7. Further flexibility in the attainable behavior of our robots can be obtained by including the actuation loading time, which results in the same DPR morphology adapting its response to can pick an object and release it at two different times as demonstrated in Movie 7.

### v3

@@ 2.2 Model derivation from mechanics-informed parameter determination
The model is further extended to capture the structure’s dynamic response, allowing us to observe the material’s viscoelastic behavior and analyze metastable cases that energy minimization alone cannot predict. The dynamic response for different unit cell geometries is shown in Figure S7b, and for different five-segmented fingers in Figure S7d, where the dynamic model predicts the same final state of the system after reaching a steady state. To explore the combined behavior of different dome geometries, we modeled a five-segment DPF topology with bistable and metastable units (see Figure 3c). We stacked four bistable units ($H_{1}=H_{2}=H_{3}=H_{4}=5$ mm) and two metastable units with the same dome height ($H_{5}=H_{6}=3$ mm). This configuration allows us to achieve one unique set point: a stable state and a time-dependent state that holds its shape for a given predictable time $\tau$ (see Figure 3c). The resulting dynamic response features five different snap-through events and two snap-backs (dome-unit resets), adequately representing the complex dynamics displayed by the DPF (Figure 3c).
2.3 Inverse co-design
@@
Based on the model’s accuracy and agreement with finite element (FE) simulations, we formulate an inverse problem to co-design diverse DPF topologies capable of reaching a desired pre-programmed set point and dynamic behavior. Since the system’s stability is governed by its stable states, the optimization algorithm searches only within feasible stable configurations (i.e., the number of active units) and their neighborhoods, resulting in high computational efficiency. To this end, we employ Bayesian optimization [51] to explore various design parameters, including the number of segments, dome height for each segment ($H_{i}$), dome thickness ($t$), unit separation for each segment ($\text{U}^{i}_{\text{sep}}$), and unit length for each segment ($\text{U}^{i}_{\text{L}}$) (See Figure 2a for parameter definitions and Figure S11 for the design space). We consider different objective functions to demonstrate the versatility of the DPF topology in encoding distinct final positions and dynamic responses with minimal actuation. First, we target a specific tip position while simultaneously maximizing stiffness (Position + Stiffness Design; see Section A.6), thereby enabling embodied control with maximum stiffness. Second, we exploit the dynamic response of the structure to encode task-specific behaviors directly into the DPF topology, enabling task planning capabilities.
Figure 4: DPF inverse design and experimental validation. (a) Predicted results for Target 1 and 3 from the optimization model (Position Design). (b) Objective function values as a function of the number of segments, with the minimum value identified for each target position. (c) The gripper achieves multiple stable kinematical states predicted by the energy-based model, which are successfully transferred to 3D-printed prototypes. (d) Stiffness determination of the DPF using a follower force approach, with Normalized strain energy vs. displacement curves, is shown for Objectives 1, 3, and 5. (e) Comparison of the proposed model and experimental results for position-only (Pos) optimization and combined position + stiffness optimization (Stf). (f) Dynamic response for a six-section DPF under two different loading profiles.
Position + stiffness design:
@@
Given a target tip coordinate [x,y] ($\text{Target}_{xy}$), the inverse problem objective function can be written as:
$\min_{H_{i},t,\text{Unit}^{i}_{L},\text{Unit}^{i}_{s}}w_{1}\left(\text{Target}_{xy}-\text{Tip}_{\text{ dis}}\right)^{2}+w_{2}\left(\frac{1}{dF_{\text{in}}/dx}\right)^{2}$ (6)
@@
where $\text{Tip}_{\text{ dis}}$ represents the tip point (see Tip in Figure 2a), which coincides with the top node of the finger’s last unit. $w_{1}$ and $w_{2}$ are the weights assigned to each objective, and $dF_{\text{in}}/dx$ is the stiffness obtained from the model (see details in section A.6.3). A domain constraint is imposed to guarantee that every dome unit is higher than the one in front ($H_{i+1}\leq H_{i}$), which in some instances can lead to up to N possible stable states (N being the number of segments). Moreover, we impose constraints for all unit separations and unit lengths to be equal for simplicity and design domain to guarantee bistability on all units, based on the phase diagram shown in Figure S1b. However, different restrictions can be imposed depending on the purpose of the DPF design, and parameters can be individually tuned to achieve a large variety of positions and stable states (see Movie 3). Our algorithm was tested by defining five different tip coordinate targets using two materials with distinct properties (NinjaFlex 85A and Cheetah 95A) optimized to achieve the desired positions. For each case, the target coordinates in $x$ and $y$ were provided to the algorithm (Target Position in Figure 4a), and the optimization was performed by incrementally adding one unit per run and calculating the objective function. To reduce the computational cost, a maximum of 8 units was specified (Figure 4b), although additional units can be included if the objective requires a larger finger geometry. The geometry with the minimum objective function value (Obj in Tables S4 and S5) was selected and plotted to verify whether the desired position was achieved (see Figure 4a). Geometric parameters, coordinate targets, and objective function values are reported in Table S4 and S5. The final stable state and its comparison with each target can be observed in Figure S12. To validate our optimization algorithm, we 3D printed the five different geometries shown in Table S9 using both tested materials (NinjaFlex 85A and Cheetah 95A) and measured the final tip position after commanded dome inversion. An average error of 7.3% was measured across all samples, showing good agreement with the optimization algorithm and the lattice model. To evaluate the impact of including stiffness in the objective function, we experimentally tested (see Figure S15) the five topologies generated via position-only optimization (Equation S.30) and position + stiffness optimization (Equation 6). Using this co-design strategy, the DPF stiffness has increased on average 1.94 times for all the test cases, achieving maximum stiffnesses of 0.55 N/mm, 0.3 N/mm, and 0.15 N/mm for cases 2, 4, and 5, respectively (see Figure 4e for case 5), which consequently can enhances the achievable grasping force.
@@
The topologies generated by this methodology exhibit deterministic behavior despite using polymeric materials, meaning that kinematic states are consistently attained as long as the gripper’s geometric parameters remain unchanged (see Figure 4c and Movie 4). Specifically, we cycle each gripper topology between the initial and fully activated states, verifying the gripper’s ability to achieve the same aperture and positioning across multiple cycles (Tests 1-3 in Figure 4c). Furthermore, we assessed material and time degradation effects by performing 300 cycles over three weeks (see Figure S16), which demonstrates that even after numerous activations and an extended period, the gripper’s performance remains within 5% of the initial design behavior. This characteristic enables robust robot design, as the bistable elements provide intrinsic kinematic control dictated by the gripper’s mechanical response. It is worth mentioning, that the pressure for activating the units is not controlled using sensors and closed-loop control, illustrating the embodied self-regulation from the designed multistability. Consequently, the position control inherent to the designed multistable structure serves as control set points that can be embodied in the robot’s morphology and achieved with a single, open-loop controlled actuator.
Dynamic response design: The unit cell morphology and co-design approach also allow us to encode desired dynamics responses in our soft robots. To this end, we utilized the model described in the previous section and added a term to account for the resetting time. Using the dynamic model (see section A.4), the metastable units’ resetting time is calculated, and the final position of the finger is determined. Given this, specific set points and dynamic behavior in the vicinity of that stable point can be designed. For this particular case, we selected a six-segment DPF with fixed dome heights for the first four units, and we utilized our optimization approach to maximize the resetting time under two different configurations. It is worth mentioning that the rest of the geometric parameters of the DPF ($U_{sep}$, $U_{L}$, $t_{lim}$) are kept constant during the optimization process, as their effect on the resetting time is negligible. At the same time, the optimization algorithm is utilized to determine the dome heights of the last two units and the remaining geometric parameters ($t$). Given this, the optimization problem can be posed as:
$\min_{H_{i},t}w_{j}\left(\frac{1}{dF_{\text{in}}/dx}\right)^{2}+w_{j+M}\frac{1}{RT_{2}-RT_{1}}$ (7)
@@
$\hskip 14.22636pt\quad H_{i+1}=H_{i},,\quad i=1,\dots,M-1$
where $M$ is the number of metastable units, $RT_{2}$ and $RT_{1}$ are the resetting time under two different pressure profiles (see Figure 4f and Figure S19a), $w_{j}$ are the optimization weights, and $G_{M}$ is the set that contains the metastable configurations. Notice that the last part of the objective function can be substituted by $w_{j+M}\left(\text{Target }RT-RT\right)$ (with Target $\text{Target }RT$ been the target resetting time) to design for a specific reset time of the units under a pressure profile. Despite setting most of the geometric parameters, we included the stiffness term in our objective function to maximize the grasping capacity for our pick-and-place application indirectly. The full dynamic response predicted by the model shows three different snap-through instability and two snap-backs in the two simulated cases (Figure 4 f). As expected, to maximize the resetting time at two different conditions, the algorithm would converge to the highest possible metastable unit (see Table S10) for the given thickness (see Figure S1b).
2.4 Dome Phalanx Robot (DPR)
To explore the capabilities of our design approach and the advantages of leveraging multistability in soft robotics, we fabricated two soft robot topologies (Dome Phalanx Robot - DPR). We demonstrate different operation modes using the proposed design methodology and FDM 3D printing to create multiple robot architectures. Specifically, we focus on three types of applications: 1) A multistable soft gripper with four different stable configurations and a metastable configuration that leverages the viscoelastic response of the material to increase the grasping force. This architecture leverages the embodied states so the gripper can perform object size and weight sorting tasks informed by its pre-programmed morphology; 2) A hexapod soft robot that walks with alternating tripod gait [8] (Dome Phalanx Walker - DPW), by generating horizontal and vertical two-way bending functions combining bistable and metastable units on different directions (see Figure 5b-d). We impose different actuator geometry and optimization constraints (detailed below) on the model for each operation mode to ensure that each actuator is accurately designed for each specific task. 3) A combination of bistable and metastable dome geometries, where the DPR’s constitutive material viscoelastic response of the material allows us to embody a pick and place robotic behavior into the geometry (see Figure 5f and Movie 6).
Figure 5: Dome Phalanx Robot (DPR) architectures and applications. a) Classification loop for the DPG. The gripper reconfigures each design’s stable state to sort the objects by size and weight (see Movie 7). b) The DPW actuator has two different zones and three bending directions. The action unit is utilized to create turning at the given pressure command. d) DPW architecture connected to create the tripod gate ($L_{1}$-$L_{4}$-$L_{5}$ and $L_{2}$-$L_{3}$-$L_{6}$). Forward units are utilized to create turns. e) Oscillating phases of the DPW with the robot at 30 psi of pressure. f) Snapshots of embodied pick-and-place tasks: A single pressure input is given for the robot to pick the object (Commanded response) and then release it depending on the viscoelastic material response (Programmed response)(see Movie 6). g) Snapshots of the walking trajectory of the DPW. A command is given by increasing the pressure in the first cycle to enable turning in both directions (Programmed response) (see Movie 9).
1) Multistable soft robot: We further leverage the characteristics of our multistable DPF by combining different static and dynamic configurations. We utilized our strategy to design a multistable Dome Phalanx Gripper - DPG with four different stable configurations and a variable dynamic behavior (See Figure  S21 a). Given this, four different object sizes are encoded into the structure, which allows for object sorting within the range of the target configurations. A metastable unit is incorporated into the tip of each DPF to enhance its classification capabilities and allow the gripper to reconfigure momentarily into a stiffer configuration, allowing it to pick objects of the same size but with different masses. The control loop is closed by combining the morphological information of the robot with a contact sensor in one of the dome units (see section A.8.1 for details) to detect when the gripper is perturbed from its design stable position. For this application, the robot base length and DPF angle with respect to the base (see Figure  S21 for reference) are incorporated, and four different apertures are programmed into the gripper morphology by minimizing for four discrete configurations (see more detail in section A.8.1). Given this, the objective function can be modified as follows:
$\min_{H_{i},t,\text{Unit}^{i}_{L},\text{Unit}^{i}_{s},\theta_{B},Base_{L}}w_{j}\left(\text{Object Size}_{j}-\text{Tip}_{\text{dis y j}}\right)^{2}+w_{M+N+j+1}\left(\frac{1}{dF_{\text{in}}/dx}\right)^{2}$ (8)
[…]
$\hskip 14.22636pt\ \quad H_{i}\in G_{M},,\quad i=M,\dots,N$
@@ 2.4 Dome Phalanx Robot (DPR)
where $N$ is the number of targeted stable states, $M$ is the number of metastable units, and $j$ is the number of test objects ($1\leq j\leq M+N$). By doing this, we can optimize for four target positions and maximum stiffness in the fully activated state (four inverted units) and one metastable unit. It is worth noting that for this case, there is a different constraint for the dome height (i.e., $H_{i+1}>H_{i}$), which allows each of the states to be useful for grasping objects while being accessible in a sequential manner (increasing the input pressure). While the model could be expanded to modify the unit cell’s dimensions further and determine the optimal number of units, we fixed the number of units at five to ensure at least one active dome per object size for simplicity. We tested our system by performing an inverse design with four different target positions (see section A.8.1 for reference), using the same methodology described previously (section 2.3). This methodology demonstrates the versatility of our DPF and provides an alternative approach to encoding multiple set points within the finger’s body by modulating its unit cell geometry. Using this approach, we embed the classification task directly into the gripper’s morphology by leveraging the encoded set points to infer object size. A contact sensor (see Section A.8.1 for details) is integrated into the first unit of one DPF, enabling the detection of perturbations from the gripper’s stable configuration. Based on this, we implement a classification loop (Figure 5a): the gripper approaches the object, evaluates contact sensor feedback, and reconfigures if the measured resistance is below a programmed threshold. Upon contact detection, the object is gently lifted with low pressurization ($P<5$ psi), and slippage is monitored by measuring the contact sensor a second time. If no slippage occurs, the object is classified into one of four size categories (see Movie 6). If the lift fails, the gripper reattempts by inverting the metastable unit at the tip, increasing stiffness to handle heavier objects of similar size (see Movie 7). This methodology enables simultaneous size and weight classification by combining the system’s mechanical response with real-time feedback from a single sensor signal. More importantly, given that the robot’s morphology purely drives the classification, there is robustness on the specific task, as it decouples from the actuation pressure.
A key characteristic of our multistable soft robots is their intrinsic capacity to attain a desired shape and exert force without the need from external work, commonly from pressure. This allows our robots to work even after sustaining significant damage, which we induce here by piercing our DPF with several needles (see Movie 7). Notably, our robot maintains its embodied capability to classify objects even under the presence of leaks (illustrated by the oscillating tufts in Movie 9). This test demonstrates the robustness of the embodied control from our approach, as well as illustrating the damage tolerance afforded by the robot’s multistability.
2) Hexapod walker: In the second application, we demonstrate the versatility of our dome unit to design a different topology, namely, soft robotic leg for a hexapod walker (see Figure 5b and 5c). The presented walker utilizes a tripod gait, which means that three legs (two on one side and one on the opposite side, i.e., $L_{1}$, $L_{5}$ and $L_{4}$ in Figure 5c) move simultaneously forward, while the other three stay in contact with the ground (See Figure 5d). Each leg is created by rotating the dome unit into different planes to create a leg that moves in two planes (see Figure 5b), which are connected by using a rigid base (see Figure S22a). The leg is divided into two zones (highlighted in purple and green in Figure 5b), which actuate in an alternating sequence to obtain the desired movement. More importantly, we leverage metastable units in the purple region to expand the range of the oscillating cycle on the phase where the legs are not touching the ground (see Figure S22e and Movie 9 for a comparison between metastable and monostable cases), utilizing the resetting time to delay the reversion of the domes while the other phase is actuated. This amplifies the overall movement, as the leg would touch the ground at a farther point than its monostable counterpart (see Figure S22e). Furthermore, the green zone is designed to move in the negative y axis (-y) and z-axis, allowing the leg to touch the ground and move backward to generate the oscillating cycle shown in Figure S22c. The zone contains three different types of units: one monostable unit (which gives the -y movement), three bistable units that provide the negative z axis (-z) movement and can reconfigure to a second stable state, and a action unit that can be either monostable or metastable. The versatility of our robots’ units allows us to have two different types of legs, front legs ($L_{1}$ and $L_{2}$) which have a metastable action unit to generate asymmetry (see Movie 9) in the movement and be able to change the walking direction (see section A.9 for details), and the middle and back legs, where we select the action units to be monostable as no asymmetry is needed. For the action unit on the front legs, the resetting time is designed to be larger than the one in the purple region, which results in a programmed response that only depends on the input pressure. As the pressure increases in one of the phases, the action unit is inverted (Commanded response in Figure 5e) inducing asymmetric locomotion in the system (moving the center of mass of the robot), which would stay active as the cycle time of each phase is lower than the resetting time of the action unit (Programmed response Figure 5e). Note that the robot can recover its forward movement by increasing the pressure on the opposite phase, restoring the symmetry in the system with the two action units inverted. Finally, the bistable units on the green zone can be inverted to morph the robot into an arch-like shape (see Figure S22b), that increases its payload capacity as it avoids the friction with the ground. As a result, the morphology of the leg embodies a mechanical logic behavior that allows the DPW to move forward, turn in both directions and reconfigure just by controlling the input pressure (see Movie 9).
@@
3) Embodied soft robotic tasks: In the second application, we utilize our ability to program our system’s dynamic response. By leveraging our dome structures’ viscoelastic response and geometry, we can pre-program the duration for which the structure remains activated before returning to its original state. Combining the meta- and bi-stability of the units enables us to design task planning operations (such as pick-and-place; see Movie 6), where the robot can pick an object by activating all units, move to a specific position, and release the object without additional interventions and with just one pressure input (i.e., eliminating the need for a release command; see Figure 5e). We tested this operation mode by utilizing the architecture designed in section 2.3 - Dynamic response design, to create a DPG architecture. The obtained results show how an object can be held while all units are inverted, even when the system releases pressure (Figure S20). However, the object is automatically released from the gripper after the metastable units reset, as desired from the programmed dynamics. Using this approach, we can program a simple pick-and-place operation by activating all domes or perform a simple pinch operation as demonstrated in Figure S20 and Movie 6. Further flexibility in the attainable behavior of our robots can be obtained by including the actuation loading time, which results in the same DPG morphology adapting its response to pick an object and release it at two different times as demonstrated in Movie 6.